# Automatically Evaluating the Paper Reviewing Capability of Large Language Models

### v2

@@ Abstract
Peer review is essential for scientific progress, but it faces challenges such as reviewer shortages and growing workloads. Although Large Language Models (LLMs) show potential for providing assistance, research has reported significant limitations in the reviews they generate. While the insights are valuable, conducting the analysis is challenging due to the considerable time and effort required, especially given the rapid pace of LLM developments. To address the challenge, we developed an automatic evaluation pipeline to assess the LLMs’ paper review capability by comparing them with expert-generated reviews. By constructing a dataset [^1] https://figshare.com/s/d5adf26c802527dd0f62 consisting of 676 OpenReview papers, we examined the agreement between LLMs and experts in their strength and weakness identifications. The results showed that LLMs lack balanced perspectives, significantly overlook novelty assessment when criticizing, and produce poor acceptance decisions. Our automated pipeline enables a scalable evaluation of LLMs’ paper review capability over time.
Automatically Evaluating the Paper Reviewing Capability of
Large Language Models
@@ 1 Introduction
Figure 1: We consider paper review task that generates a summary of paper, strengths, weaknesses, and the final judgement. Our goal is to examine the level of agreement between LLMs and human experts in reviewing papers, based on their feedback targets and aspects.
@@
Reviewing academic papers lies at the heart of scientific advancement, but it demands substantial expertise, time, and effort. The peer review system faces several challenges, including a growing number of submissions that outpace the reviewer availability, lack of incentives, and reviewer fatigue (Tropini et al., 2023; Horta and Jung, 2024; Hossain et al., 2025). While Large Language Models (LLMs) hold the potential to assist reviewers by reviewing papers automatically (Hosseini and Horbach, 2023; Robertson, 2023), prior research has reported significant limitations in their performance (Du et al., 2024; Liang et al., 2024; Zhou et al., 2024). For example, studies have highlighted that LLM-generated reviews often lack accuracy, detail, and specificity when compared to reviews written by human experts (Zhou et al., 2024; Mostafapour et al., 2024).
While these findings are informative, they do not sufficiently clarify the precise differences between expert-generated reviews and LLM-generated reviews. Specifically, it is still unclear to what extent LLMs excel or fall short in different aspects of reviewing compared to human experts. Addressing the question requires systematic quantitative analysis of review data, but such analysis is challenging to scale due to the significant time and effort required from researchers.
To address this gap, we introduce an automated pipeline designed to systematically analyze these differences. Our approach is to automatically annotate the strengths and weaknesses identified in reviews based on targets (e.g., problem, methodology, and experiment) and their associated aspects (e.g., validity, clarity, and novelty) and examine the agreement between LLMs and human experts (Figure 1). By introducing a systematic framework for examining the strengths and limitations of LLMs in academic review, this work offers valuable insights into improving their performance and enhancing their potential role in assisting the review process.
Figure 2: The overall evaluation process. Given a paper, we extract strengths and weaknesses from review data on the OpenReview platform. To identify key strengths and weaknesses that influence the final acceptance, we extracted them from the meta-review and augmented details from reviewer comments to make them self-contained. We then compared these with LLM-identified strengths and weaknesses, based on their feedback target and aspect. The evaluation is conducted automatically, enabling a scalable and longitudinal evaluation over time.
Our study leverages a dataset of 676 papers and their review data that has been collected from OpenReview [^2] https://openreview.net/ for ICLR conferences spanning 2021 to 2024. We extracted the strengths and weaknesses highlighted in the meta-reviews using gpt-4o333gpt-4o-2024-08-06, which will be compared with LLM-identified strengths and weaknesses. Then, we developed a coding schema (Table 1) for annotating the targets and aspects of these strengths and weaknesses, by surveying 9 AI paper submission guidelines and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). Based on the schema, we manually annotated 327 strengths and weaknesses of 68 randomly sampled papers, providing a basis for building an automatic annotation tool. Our LLM-powered automatic annotation tool achieved 0.85 (target) and 0.86 (aspect) inter-rater reliability (IRR) with human-annotated results, showing high level of consistency and accuracy. Overall, we identified 3,657 review items (1,231 strengths and 2,416 weaknesses) from the review dataset.
We evaluated 8 LLMs (4 GPT, 2 Llama, and 2 DeepSeek family) for their paper review capability. After generating reviews for each of the 676 papers, we analyzed the agreement between LLMs’ and experts’ reviews based on their target and aspect assigned to strengths and weaknesses. The results showed that: 1) LLMs lack balanced perspective compared to human experts, 2) LLMs significantly neglect novelty assessment for evaluating papers’ weaknesses, and 3) the paper acceptance decisions are not accurate. The findings are consistent for all the LLMs, highlighting clear opportunities for improving their reasoning capability.
@@
We release a dataset comprising 68 papers, experts reviews, 3,657 strengths and weaknesses identified from the reviews with automatically annotated targets and aspects, LLM-generated reviews from 8 LLMs, and a total of 43,042 strengths and weaknesses identified from the LLMs with their annotated targets and aspects.
2 Task
We define the paper review generation task as follows: given a research paper, 1) summarize main points, 2) identify a list of strengths and weaknesses, and 3) predict the final acceptance of the paper. This task offers direct value to various user groups (e.g., authors who want to get initial feedback on their draft, or reviewers who want to examine diverse view points) by providing actionable feedback for improving their papers. While valuable, evaluating papers based on research standards (e.g., novelty, rigor, and clarity) is difficult for LLMs as it requires significant expertise.
@@ 3.1 Collecting Review Data
3.2 Identifying Strengths and Weaknesses
@@
One of the challenges in identifying the strengths and weaknesses of these papers is that each review consists of multiple blocks, including a meta-review and individual review texts from several reviewers. To address the challenge, our approach is to use meta-review, a final review from a qualified expert that summarizes reviews and highlights important strengths and weaknesses for supporting the final decision. As the meta-review does not capture all the details, we created self-contained strengths and weaknesses by 1) extracting them from the meta-review and 2) augmenting these extracted elements with detailed comments from individual reviews (non-meta). We designed a prompting chain that consists of three prompts (Appendix A.1.1). After excluding withdrawn submissions that lack meta-reviews, 14,922 submissions remained.
Target
Code Definition (The review addresses …)
Problem Motivation, task definitions, and problem statements.
Method Proposed approach, techniques, algorithms, or datasets.
Experiment Experimental setup, results, and analysis.
Conclusion Findings, implications, discussions, and takeaways.
Aspect
Code Definition (The review addresses …)
Impact Significance or practical influence of the work.
Novelty Originality of the submission compared to prior research.
Clarity Readability, ambiguity, or communication aspects.
Validity Soundness, completeness, and rigor.
Table 1: The coding schema. To identify codes for targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated), we surveyed 9 AI paper submission guidelines (Appendix  A.2.1) and prior research on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022).
4 Building an Automatic Annotator Based on Target and Aspect
The central goal of this paper is to analyze where LLMs excel and fall short in reviewing papers, compared to human experts. To achieve the goal, we 1) annotate each of the strengths and weaknesses identified by LLMs and experts and 2) examine the agreement between them based on the annotation results. The analysis offers insights into the distinct contributions and limitations of LLMs in reviewing papers, informing strategies to foster more effective human-LLM collaboration in reviewing papers.
4.1 Developing a Coding Scheme
@@
Our focus in classifying strengths and weaknesses lies in two key dimensions: targets (i.e., what the review praises or critiques) and aspects (i.e., the specific elements of the target being evaluated). To build an initial codebook, we surveyed 9 AI paper submission guidelines (Appendix A.2.1) and extracted target-aspect pairs from each statement in the guidelines (e.g., “The paper should state the full set of assumptions of all theoretical results if the paper includes theoretical results.” yields the target Theory and aspect Completeness). We also reviewed related work on the analysis of paper review data (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). After identifying 33 targets and 13 aspects, we merged similar items to create simple and distinct categories, resulting in 7 targets and 4 aspects. Table 1 shows our final coding scheme.
4.2 Building an Automatic Classifier Based on Human Annotation
@@
Based on the coding scheme, we annotated targets and aspects of strengths and weaknesses to produce ground truth for developing an automatic annotator. We randomly sampled 68 papers from our review dataset, yielding 327 instances of strengths and weaknesses. Two authors annotated each instance together, resolving any conflicts. Most conflicts arose when an instance illustrated multiple points. For example, an instance such as “**Technically sound with a strong foundation**: The paper’s technical foundation is evident in its bi-level optimization framework, … Technical novelty also arises from using supermartingale constraints on the barrier function …” could correspond to both Validity and Novelty aspect. Two authors finalized the annotation through discussions, focusing on the main point or root cause of the issue. In the example, we annotated Validity, as the strength mainly praises the technical soundness, as shown in the header.
We then designed prompts to automatically annotate the instances, assigning a target and aspect label to each. Specifically, we designed four prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness A.2.2. Table 2 shows the inter-rater reliability (IRR) between author annotations and LLM annotations. Classification using o3-mini achieved the IRR scores of 0.85 for targets and 0.86 for aspects. Given the high IRR and its relatively low computational cost, we used o3-mini for the automatic annotation of both target and aspect in the main evaluation. Moreover, an examination of the confusion matrix (Appendix A.2.3) suggests that the errors tend to occur in semantically related categories, indicating that the misclassifications are not arbitrary but rather reflect subtle ambiguities inherent in the data.
@@
gpt-4o-mini 0.75 0.80
gpt-4o 0.87 0.83
o3-mini 0.85 0.86
Table 2: Inter-Rater Reliability between annotations of authors and LLMs for the target and aspect.
@@ 5 Evaluation
Overall Strength Weakness
Model F1 Precision Recall F1 Precision Recall F1 Precision Recall
DeepSeek-R1 0.373 0.314 0.460 0.341 0.254 0.520 0.400 0.379 0.424
o1-mini 0.359 0.283 0.491 0.331 0.232 0.578 0.385 0.343 0.439
o1 0.355 0.300 0.436 0.318 0.234 0.495 0.388 0.377 0.400
DeepSeek-V3 0.351 0.300 0.421 0.330 0.246 0.501 0.368 0.362 0.374
Llama-405B 0.350 0.323 0.381 0.349 0.279 0.465 0.350 0.371 0.331
gpt-4o 0.349 0.287 0.442 0.342 0.252 0.533 0.354 0.325 0.388
gpt-4o-mini 0.344 0.289 0.427 0.335 0.246 0.522 0.353 0.337 0.369
Llama-70B 0.339 0.302 0.388 0.338 0.260 0.481 0.341 0.350 0.332
Table 3: Overall performance of alignments on strengths and weaknesses between experts-identified and LLM-identified reviews. The metrics were computed by comparing the (target, aspect) set between experts’ and LLMs’ review. DeepSeek-R1 achieved the best performance, o1-mini achieved superior recall, and Llama-405B achieved superior precision, compared to other models.
The goal of our evaluation is to analyze the strengths and weaknesses of a given paper as identified by LLMs, comparing them with those identified by human experts. Note that our evaluation does not consider the correctness of the identified strengths and weaknesses because our focus is comparing perspectives in reviewing papers for both groups, not the content itself.

### v3

# Mind the Blind Spots: A Focus-Level Evaluation Framework for LLM Reviews

@@ Abstract
Peer review underpins scientific progress, but it is increasingly strained by reviewer shortages and growing workloads. Large Language Models (LLMs) can automatically draft reviews now, but determining whether LLM-generated reviews are trustworthy requires systematic evaluation. Researchers have evaluated LLM reviews at either surface-level (e.g., BLEU and ROUGE) or content-level (e.g., specificity and factual accuracy). Yet it remains uncertain whether LLM-generated reviews attend to the same critical facets that human experts weigh—the strengths and weaknesses that ultimately drive an accept-or-reject decision. We introduce a focus-level evaluation framework that operationalizes the focus as a normalized distribution of attention across predefined facets in paper reviews. Based on the framework, we developed an automatic focus-level evaluation pipeline based on two sets of facets: target (e.g., problem, method, and experiment) and aspect (e.g., validity, clarity, and novelty), leveraging 676 paper reviews [^1] https://figshare.com/s/d5adf26c802527dd0f62 from OpenReview that consists of 3,657 strengths and weaknesses identified from human experts. The comparison of focus distributions between LLMs and human experts showed that the off-the-shelf LLMs consistently have a more biased focus towards examining technical validity while significantly overlooking novelty assessment when criticizing papers.
Mind the Blind Spots:
A Focus-Level Evaluation Framework for LLM Reviews
@@ 1 Introduction
Figure 1: We introduce a focus-level evaluation framework for assessing LLM reviews, which computes focus distributions and compares them against human reviews based on predefined facets. The focus-level evaluation offers actionable insights into how to improve LLMs’ paper review capability and how to most effectively leverage LLM reviews in the peer review process.
@@
Reviewing academic papers lies at the heart of scientific advancement, but it requires substantial expertise, time, and effort. The peer review system faces several challenges, including a growing number of submissions that outpace the reviewer availability, lack of incentives, and reviewer fatigue (Tropini et al., 2023; Horta and Jung, 2024; Hossain et al., 2025). Large Language Models (LLMs) hold the potential to assist the peer review process by automatically reviewing papers (Hosseini and Horbach, 2023; Robertson, 2023), but can we trust LLM-generated reviews? Evaluating the quality of reviews is inherently complex due to their multi-dimensional nature. Researchers have employed various metrics for the evaluation such as surface-level (e.g. linguistic similarity to human reviews), content-level (e.g., relevance, specificity, and factual accuracy), and decision-level (e.g., accept/reject classification accuracy) metrics (Ramachandran et al., 2017; Du et al., 2024; Liang et al., 2024; Zhou et al., 2024).
However, existing evaluations fail to assess whether LLM reviews comprehensively address critical dimensions of papers. Evaluating the focus of reviews is crucial because reviews with poor focus can negatively impact reviewers, even if they are accurate, relevant, and specific. For example, reviews that overly concentrate on methodological details while completely neglecting the novelty aspect of the proposed method could fail to suggest meaningful feedback, diverging from how expert reviewers assess the submission. It could also mislead junior reviewers by promoting incomplete perspectives and reinforce shallow paper review practices. Despite such importance, few attempts have been made to systematically evaluate whether the focus of LLM reviews aligns with that of expert reviews. Conducting the focus-level evaluation of LLM reviews is useful to reveal the blind spots of LLM reviews along with their central focus, offering important insights into how human reviewers can most effectively leverage LLM reviews in the peer review process. Moreover, it provides a concrete foundation for guiding LLM training toward more balanced and expert-aligned review behavior.
We introduce a framework for focus-level evaluation of LLM reviews, which systematically analyzes where the reviews direct their praise and criticism based on facets considered important in peer review (Figure 1). Given an LLM, the framework computes a focus distribution, a normalized distribution of how frequently review points (e.g., a list of strengths and weaknesses) address predefined facets (e.g., problem, method, and experiments) by leveraging a paper review dataset. The focus distribution can be computed by an automatic annotator that assigns a facet for each review point, enabling a fully automatic evaluation. The interpretable nature of the focus distribution provides actionable insights by clearly revealing which facets LLMs tend to emphasize or overlook in comparison to human experts.
To apply this framework for analyzing LLM-generated reviews in the context of AI conferences, we implemented a focus-level evaluation pipeline (Figure 2). We identified the facets that constitute review focus, by surveying 9 paper submission guidelines from AI conferences and prior literature on review analysis (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). We define two sets of facets: target (what review points praise and critique such as problem, method, and experiment) and aspect (which criteria is being evaluated such as validity, clarity, and novelty), which are key elements in analyzing paper reviews (Ghosal et al., 2022; Lu et al., 2025). We identified 7 facets for the target and 5 facets for the aspect (Table 1). Next, we developed an automatic annotator for computing the focus distributions based on the target and aspect, which assigns a target and aspect label for a strength and weakness point in a review. The annotator showed substantial agreement with human annotators, achieving IRR (Cohen’s kappa (Cohen, 1960)) of 0.81 for target and 0.79 for aspect.
As a benchmark dataset for our focus-level evaluation pipeline, we constructed a dataset of 676 papers and their review data from OpenReview for ICLR conferences spanning 2021 to 2024. Then we computed and compared the focus distributions of human and LLM reviews using the evaluation pipeline (Figure 4), and we also measured text similarities between the reviews. Specifically, we evaluated 8 LLMs (4 GPT, 2 Llama, and 2 DeepSeek family) to analyze their review focus. We also evaluated MARG (D’Arcy et al., 2024) as a novel review generation technique and a fine-tuned gpt-4o using our dataset. The results showed that:
  • LLMs struggle to identify key targets and aspects in their reviews. Even the top-performing model reached an F1 score of 0.373 when matching human reviewers on the targets and aspects in each review point.
  • LLMs’ review focus was biased towards examining technical validity, consistently overlooking novelty assessment in weaknesses – a critical limitation in paper review.
  • The fine-tuned model produced focus distributions most closely aligned with that of humans, compared to models using prompting alone.
  • The models demonstrated strengths in distinct areas. While the fine-tuned model produced the closest focus distributions, Llama-405B achieved the highest text similarity. It highlights the importance of holistic evaluation to capture the diverse aspects of review quality.
Figure 2: The overall process of automated focus-level evaluation. We first extracted strengths and weaknesses from review data on the OpenReview platform as the expert reviews. To identify key strengths and weaknesses influencing the final acceptance, we extracted them from the meta-review and augmented details from individual reviewer comments. Each strength and weakness was then annotated with a target and aspect by our automatic annotator. Finally, we computed the focus distributions by normalizing the frequency of annotated targets and aspects, and compare this distribution with that of LLM reviews.
@@
We release a dataset comprising 676 papers, expert reviews, 3,657 strengths and weaknesses identified from the expert reviews with automatically annotated targets and aspects, LLM-generated reviews from 8 LLMs, and a total of 43,042 strengths and weaknesses extracted from the LLMs, each annotated with corresponding targets and aspects.
2 A Framework for Focus-Level Evaluation of LLM Reviews
We propose a focus-level evaluation framework to systematically analyze what aspects LLMs emphasize or overlook when reviewing scientific papers. To enable interpretable and automated assessments of LLM behavior in reviewing, we aim to reveal the distribution of attention an LLM allocates to different review facets when identifying strengths and weaknesses in submissions. Specifically, we define a focus of the review to be compared for focus-level evaluation as follows:
Let (i) $L$ be an LLM, (ii) $A=\{a_{1},a_{2},...,a_{N}\}$ be a list of facets where each facet denotes a distinct criteria (e.g., problem, method, and experiment), and (iii) $P=\{p_{1},p_{2},...,p_{M}\}$ be a corpus of paper submissions. The focus-level evaluation $E(L,A,P)$ produces two focus distributions $F^{+}$ and $F^{-}$ where $F^{+}$ denotes the distribution when identifying strengths of the submissions and $F^{-}$ for weaknesses. The focus distribution $F=(f_{1},f_{2},...,f_{N})$ can be represented as a normalized vector where $f_{i}$ denotes the relative frequency of review points (i.e., strengths for $F^{+}$ and weaknesses for $F^{-}$) that discuss the facet $a_{i}$, when $L$ generates reviews for paper submissions in $P$.
To assess LLM behavior, our framework compares focus distributions with those from human expert reviews. Researchers can specify the set of facets $A$ and the paper corpus $P$ based on the goals of their analysis, allowing flexible and targeted focus-evaluation.
Based on this framework, we implement an automatic focus-level evaluation pipeline to understand LLM’s behavior in reviewing AI papers. Figure 2 illustrates the process of our focus-level evaluation pipeline. Our approach consists of three steps. (i) Collect an expert review dataset from ICLR conferences and extract strengths and weaknesses of the submissions for computing focus distributions of human experts (Section 3), (ii) Define facets based on paper submission guidelines of AI conferences and build an automatic annotator based on the facets (Section 4), and (iii) Compute and analyze the focus distributions of LLMs and human experts in reviewing AI papers (Section 5).
@@ 3 Constructing Expert Review Dataset
The focus-level evaluation framework requires a corpus of paper submissions $P$. We collected the review data from OpenReview platform and extracted the strengths and weaknesses of papers for computing focus distributions of human experts.
@@ 3.1 Collecting Review Data
3.2 Extracting Strengths and Weaknesses
@@
One of the challenges in identifying the strengths and weaknesses of these papers is that each review consists of multiple blocks, including a meta-review and individual reviews from several reviewers. To address the challenge, our approach is to use a meta-review, a final review from a qualified expert that summarizes reviews and highlights important strengths and weaknesses for supporting the final decision. As the meta-review does not capture all the details, we created self-contained strengths and weaknesses by 1) extracting them from the meta-review and 2) augmenting these extracted elements with detailed comments from individual reviews (non-meta). We designed a prompting chain that consists of three prompts (Appendix A.1.1).
4 Developing an Automatic Focus-level Evaluation Method
To enable a fully automated evaluation using the proposed focus-level evaluation framework, we first define a set of facets and then develop an automatic annotator. We then compute focus distributions based on the annotated facets to analyze how LLMs and human reviewers differ in their focus of reviewing.
4.1 Defining Facets from Guidelines
@@
To build an initial set of facets, we surveyed 9 AI paper submission guidelines (Appendix A.2.1) and extracted target-aspect pairs from each statement in the guidelines (e.g., “The paper should state the full set of assumptions of all theoretical results if the paper includes theoretical results.” yields the target Theory and aspect Completeness). To ensure comprehensive coverage of facets, we also reviewed literature that analyzes paper review data (Chakraborty et al., 2020; Ghosal et al., 2022; Yuan et al., 2022). After identifying 33 targets and 13 aspects, we merged similar items to create simple and distinct categories, resulting in 7 targets and 4 aspects (Table 1). The definition of each target and aspect facet is available in Appendix A.2.2.
Target Aspect
Problem Impact
Prior Research Novelty
Method Clarity
Theory Validity
Experiment Not-specific
Conclusion
Paper
Table 1: Our research focuses on two sets of facets: target and aspect. Detailed definitions of the facets are available in Appendix A.2.2.
4.2 Building Automatic Annotators
@@
Based on the identified facets, we annotated targets and aspects of strengths and weaknesses to produce ground truth for developing an automatic annotator. We randomly sampled 68 papers from our review dataset, yielding 327 instances of strengths and weaknesses. Two authors — one author is experienced in qualitative research in HCI and the other author has prior publications in the field of AI/NLP — synchronously decided each label together, resolving any conflicts. Most conflicts arose when an instance illustrated multiple points. For example, an instance such as “**Technically sound with a strong foundation**: The paper’s technical foundation is evident … Technical novelty also arises from using supermartingale constraints …” could correspond to both Validity and Novelty aspect. Two authors finalized the annotation through discussions, focusing on the main point or root cause of the issue. In the example, we annotated Validity, as the strength mainly praises the technical soundness, as shown in the header wrapped in “**”.
@@
gpt-4o-mini 0.69 0.71
gpt-4o 0.83 0.75
o3-mini 0.81 0.79
Table 2: Inter-Rater Reliability (Cohen’s kappa Cohen (1960)) between annotations of authors and LLMs.
We then designed prompts to automatically annotate the instances, assigning a target and aspect label to each. Specifically, we designed four prompts where each corresponds to one of the four combinations of target/aspect and strength/weakness A.2.3. Table 2 shows the Inter-Rater Reliability (IRR, Cohen’s kappa Cohen (1960)) between human and LLM annotations for three language models. Annotation using o3-mini achieved the IRR scores of 0.81 for targets and 0.79 for aspects, indicating substantial agreement Cohen (1960). Given the high IRR and its relatively low computational cost compared to other two models, we used o3-mini for the automatic annotation of both target and aspect in the main evaluation. Moreover, an examination of the confusion matrix (Appendix A.2.4) suggests that the errors tend to occur in semantically related categories, indicating that the misclassifications are not arbitrary but rather reflect subtle ambiguities inherent in the data.
4.3 Computing Focus Distributions
Building on the defined facets and the automatic annotation method, we assign a target and aspect label to each strength and weakness point, using the automatic annotator. We then compute the normalized frequency of these labels to derive focus distributions of targets and aspects, respectively. Separate distributions are calculated for strengths and weaknesses, resulting in four distinct focus distributions. These focus distributions illustrate how LLMs and human reviewers allocate their attention across the different facets of a paper.
@@ 5 Evaluation
Focus similarity Text similarity
Model KL Divergence Overall F1 Strength F1 Weakness F1 ROUGE-L BERTScore BLEU-4
gpt-4o-mini 0.081 0.344 0.335 0.353 0.197 0.883 0.076
gpt-4o 0.082 0.348 0.342 0.354 0.202 0.885 0.079
o1-mini 0.090 0.359 0.331 0.385 0.179 0.878 0.059
o1 0.097 0.355 0.318 0.388 0.170 0.869 0.032
DeepSeek-R1 0.120 0.373 0.341 0.400 0.156 0.874 0.045
Llama-70B 0.136 0.339 0.338 0.341 0.215 0.882 0.076
Llama-405B 0.145 0.349 0.349 0.350 0.218 0.884 0.089
DeepSeek-V3 0.151 0.350 0.330 0.368 0.199 0.880 0.069
gpt-4o (FT) 0.022 0.306 0.280 0.322 0.194 0.882 0.081
MARG 0.113 0.346 0.346 0.160 0.854 0.011
Table 3: Overall performance by comparing expert reviews and LLM reviews. For focus similarity, we computed an average of the KL divergences of four focus distributions (strength/target, weakness/target, strength/aspect, and weakness/aspect) between LLM and expert reviews. The overall, strength, and weakness F1 scores were computed by comparing the (target, aspect) set between expert and LLM reviews. The text similariy metrics were computed between LLM reviews and expert reviews. The results highlight different areas of excellence across models (gpt-4o (FT): the highest focus distribution similarity, DeepSeek-R1: the best agreement on (target, aspect) labels, Llama-405B: the highest text similarity score.)
5.1 Setup
Data.